OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (3, 1) and (2, 9)? y=1/6x+0.5 is this right?

OpenStudy (owlcoffee):

try replacing the points on the line equation to verify if the points are contained in the line.

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (3,1)...x1 = 3 and y1 = 1 (2,9)...x2 = 2 and y2 = 9 now sub slope(m) = (9 - 1) / (2 - 3) slope(m) = 8/-1 or just -8 now use y = mx + b form slope(m) = -8 use either of your points...(3,1)...x = 3 and y = 1 now we sub and solve for b, the y intercept y = mx + b 1 = -8(3) + b 1 = -24 + b 1 + 24 = b 25 = b so your slope intercept line is : y = -8x + 25 check.. y = -8x + 25 (3,1)....x = 3 and y = 1 1 = -8(3) + 25 1 = -24 + 25 1 = 1 (correct)

OpenStudy (texaschic101):

any questions at all ?

OpenStudy (anonymous):

so the answer is y=8x+25?

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

I'll save the example that you gave me to help me on other problems.